Jocelynn:

It cost Evan $17.70 to send 177 text messages. How many text messages did he send if he spent $19.10?

AnimeLOL:

ill help :)

HeartTricker:

Welcome, so 17.70 is \[177\times 10\] what would 19.10 equal?

Jocelynn:

@hearttricker wrote:
Welcome, so 17.70 is \[177\times 10\] what would 19.10 equal?
ion know its hard for me to pay attention so am not sure

snowflake0531:

We can set this up as a proportion 17.7 -> 177 19.1 -? ? \[\frac{ 17.7 }{ 177 } = \frac{ 19.1 }{ x }\] cross multiply \[17.7x=19.1(177)\] can you solve from there?